# The $u$-invariant of function fields in one variable

### v1

@@ 5. The generalised strong $u$-invariant
We are now prepared to study the $u$-invariant of a function fields in one variable. Let $K$ be a field with $\mathsf{char}(K)\neq 2$. As in [4, Sect. 5], we define
@@
We call this the generalised strong $u$-invariant of $K$. For nonreal fields, $\widehat{u}$ coincides with the ‘strong $u$-invariant’ defined in [15].
@@
We have $u(K(X))\geqslant 2\,\mathsf{sup}\{u(L)\mid L/K\mbox{ finite field extension}\}$.
@@
See [5, Proposition 6.1]. ∎
@@
The field $K$ is called quadratically closed if ${K}^{\times}={K}^{\times 2}$, and it is called hereditarily quadratically closed if ${L}^{\times}={L}^{\times 2}$ for every finite field extension $L/K$.
@@
If $K(\sqrt{-1})$ is hereditarily quadratically closed, then $\widehat{u}(K)=1$, and otherwise $\widehat{u}(K)\geqslant 2$.
@@
If $K(\sqrt{-1})$ is hereditarily quadratically closed, then $u(F)=2$ for every function field in one variable $F/K$, by [12, Theorem], so in particular $\widehat{u}(K)=1$. Assume now $K(\sqrt{-1})$ is not hereditarily quadratically closed. Hence we may take a finite field extension $K^{\prime}/K(\sqrt{-1})$ such that $K^{\prime}$ is not quadratically closed. Then $u(K^{\prime})\geqslant 2$. As $K^{\prime}(X)$ is a function field in one variable over $K$ and $u(K^{\prime}(X)))\geqslant 2u(K^{\prime})\geqslant 4$, by 5.1, we conclude that $\widehat{u}(K)\geqslant 2$. ∎
Note that by its definition, $\widehat{u}(K)$ is always in $\frac{1}{2}\mathbb{N}\cup\{\infty\}$, but we do not know whether it is always in $\mathbb{N}$. We do not even have any example where $\widehat{u}(K)$ would be not a power of $2$.
@@
Let $v$ be a valuation on $K$ with $\mathsf{char}(Kv)\neq 2$. Then
@@
[…]
$[vK:2vK]\cdot u(E)=[wF:2wF]\cdot u(Fw)\leqslant u(F)\leqslant 2\widehat{u}(K)\,.$
@@
By 5.1, we have that $2u(K^{\prime})\leqslant u(K^{\prime}(X))\leqslant 2\widehat{u}(K^{\prime})$. Hence $u(K^{\prime})\leqslant\widehat{u}(K^{\prime})$.
Consider a function field in one variable $F^{\prime}/K^{\prime}$. Consider an anisotropic torsion form $\varphi$ over $F^{\prime}$. There exists a finitely generated subextension $F/K$ of $F^{\prime}/K$ such that $\varphi$ is defined over $F$ and is a torsion form over $F$. Moreover, we may choose $F$ to be transcendental over $K$. Then $F/K$ is a function field in one variable and $\varphi$ is an anisotropic torsion form over $F$. Hence $\mathsf{dim}(\varphi)\leqslant u(F)\leqslant 2\widehat{u}(K)$. Since this holds for any anisotropic torsion form $\varphi$ over $F^{\prime}$, it follows that $u(F^{\prime})\leqslant 2\widehat{u}(K)$. Having this for any function field in one variable $F^{\prime}/K^{\prime}$, we conclude that $\widehat{u}(K^{\prime})\leqslant\widehat{u}(K)$.
Assume now that $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree. It remains to show that in this case $\widehat{u}(K)\leqslant\widehat{u}(K^{\prime})$. Consider a function field in one variable $F/K$ and an anisotropic torsion form $\varphi$ over $F$. There exists a function field in one variable $F^{\prime}/K^{\prime}$ such that $F\subseteq F^{\prime}$ and $F^{\prime}=K^{\prime}F$. Since $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree, so is $F^{\prime}/F$. Therefore, Springer’s Theorem for odd degree extensions [17, Chap. VII, Theorem 2.7] yields that $\varphi$ remains anisotropic over $F^{\prime}$. Since it also remains a torsion form over $F^{\prime}$, we obtain that $\mathsf{dim}(\varphi)\leqslant u(F^{\prime})\leqslant 2\widehat{u}(K^{\prime})$. Since this holds for all anisotropic torsion forms $\varphi$ over $F$, we see that $u(F)\leqslant 2\widehat{u}(K^{\prime})$. Having this for all function fields in one variable $F/K$, we conclude that $\widehat{u}(K)\leqslant\widehat{u}(K^{\prime})$. ∎
Note in the context of 5.5 that, for a general algebraic extension $K^{\prime}/K$, one may have that $\widehat{u}(K^{\prime})<\widehat{u}(K)$: in view of 5.3, it suffices to have that $K(\sqrt{-1})$ is not hereditarily quadratically closed and to take for $K^{\prime}$ an algebraic closure of $K$.
Question 5.6.
@@
Assume that $u(K(X,Y))=4u(K)$. Then $\widehat{u}(K)=u(K)$.
@@
By 5.5, we have $u(K)\leqslant\widehat{u}(K)$. To show the converse inequality, we need to show that $u(F)\leqslant 2u(K)$ for an arbitrary function field in one variable $F/K$. Given such $F/K$, there exist $x,y\in F$ such that $F/K(x,y)$ is a finite purely inseparable field extension. In particular, $K(x,y)$ contains a subfield $F_{0}$ which is isomorphic to $F$ and such that $F/F_{0}$ is a direct limit of odd degree extensions, and we obtain by using Springer’s Theorem for odd degree extensions [17, Chap. VII, Theorem 2.7] that $u(F)=u(F^{\prime})\leqslant u(K(x,y))\leqslant u(F)$, whereby $u(F)=u(K(x,y))$. We may thus assume that $F=K(x,y)$ for some $x,y\in F$. Then $F$ is the residue field of a discrete valuation on $K(X,Y)$. It follows by 4.1 and the hypothesis that $4u(K)=u(K(X,Y))\geqslant 2u(F)$, whereby $u(F)\leqslant 2u(K)$, as desired. ∎
By 5.1, we have $u(K(X))\geqslant 2u(K)$. This has an analogue for the strong $u$-invariant.
@@
Consider a function field in one variable $F/K$. Then $F(X)/K(X)$ is function field in one variable and $u(F(X))\geqslant 2u(F)$, by 5.1. Therefore we have $2\widehat{u}(K(X))\geqslant u(F(X))\geqslant 2u(F)$. Having this for every function field in one variable $F/K$ yields the claimed inequality. ∎
Question 5.9.
@@
Let $K$ denote the quadratic closure of $\mathbb{Q}$. Choose an arbitrary proper finite field extension $L/K$ (e.g. $L=K(\sqrt[5]{6})$). Then $L$ is nonreal, and by [17, Chap. VII, Cor. 7.11] we have $|{L}^{\times}/{L}^{\times 2}|=\infty$. In particular $u(L)\geqslant 2$. It follows by 5.1 that $u(K(X))\geqslant 2u(L)\geqslant 4>2=2u(K)$.
@@ 6. Function fields in one variable over a henselian valued field
We will now prove our main result and subsequently present various situations where it can be applied. We need a last step of preparation. For a rank-$1$ valuation $w$ on a field $F$, we denote by $F^{w}$ the corresponding completion. Let $K$ be a field with $\mathsf{char}(K)\neq 2$.
@@
Let $v$ be a henselian rank-$1$ valuation on $K$ and $\mathsf{char}(Kv)\neq 2$. Let ${F/K}$ be a function field in one variable and $w$ a rank-$1$ valuation on $F$ with $\mathcal{O}_{v}\subseteq\mathcal{O}_{w}$. Then
@@
Since $2\in{\mathcal{O}}^{\times}_{v}\subseteq{\mathcal{O}}^{\times}_{w}$, we have that $\mathsf{char}(Fw)\neq 2$. Note that $w$ extends naturally to a complete valuation $w^{\prime}$ on $F^{w}$ with $w^{\prime}F^{w}=wF$ and $F^{w}w^{\prime}=Fw$. Hence $w^{\prime}$ is a complete rank-$1$ valuation on $F$ and $\mathsf{char}(F^{w}w^{\prime})\neq 2$. It follows by [13, Theorem 1.3.1 (Hensel’s Lemma)] that $1+\mathfrak{m}_{w^{\prime}}\subseteq{(F^{w})}^{\times 2}$. The equality on the left follows now from 4.2 applied to $(F^{w},w^{\prime})$. It remains to show that $[wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$.
@@
$(1)$ Suppose that $Fw/Kv$ is a finite field extension and $wF/vK$ is a finitely generated $\mathbb{Z}$-module with $\mathsf{rrk}(Fw/Kv)=1$. It follows by 2.1 that $[wF:2wF]=2\cdot[vK:2vK]$. By 5.5 we have that $u(Fw)\leqslant\widehat{u}(Kv)$. From this the claimed inequality follows.
[…]
$(3)$ Suppose that $Fw/Kv$ is an algebraic field extension and $wF/vK$ is a torsion group. Then $u(Fw)\leqslant\widehat{u}(Kv)$ and $[wF:2wF]\leqslant[vK:2vK]$ by 2.1. In this case we obtain even that $[wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$.
$(4)$ Assume finally that $w|_{K}$ is trivial. In this case $w$ is a discrete valuation, so we may assume that $wF=\mathbb{Z}$, and the residue field $Fw$ is a finite extension of $K$. Hence $v$ extends to a valuation $v^{\prime}$ on $Fw$, and as $v$ is henselian, so is $v^{\prime}$. It follows by 4.2 that $u(F^{w})=[wF:2wF]\cdot u(Fw)=2u(Fw)$ and $u(Fw)=[v^{\prime}(Fw):2v^{\prime}(Fw)]\cdot u((Fw)v^{\prime})$. Since the extension $(Fw)v^{\prime}/Kv$ is finite, we have that $u((Fw)v^{\prime})\leqslant\widehat{u}(Kv)$, by 5.5. Furthermore, $[v^{\prime}(Fw):vK]\leqslant[Fw:K]<\infty$, whence $[v^{\prime}(Fw):2v^{\prime}(Fw)]=[vK:2vK]$ by 2.1. Thus $u(Fw)=[vK:2vK]\cdot u((Fw)v^{\prime})\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. ∎
We come to our generalisation of [28, Theorem 3] and [4, Cor. 6.4]. The proof crucially depends on the local-global principle for isotropy from [23, Cor. 3.9], or more precisely, on the slightly extended version provided in [3, Theorem 4.4], where the valuation on the base field is assumed of rank $1$ and henselian, but not necessarily complete.
@@
By 5.4, we have $\widehat{u}(K)\geqslant[vK:2vK]\cdot\widehat{u}(Kv)$, so it remains to show the converse inequality.
@@
[…]
@@
We set $K_{0}=\mathbb{F}(S\cap K)$. Since the absolute transcendence degree of $K_{0}$ is bounded by $|S|$, it follows by [13, Theorem 3.4.3] that $\mathsf{rk}(v|_{K_{0}})\leqslant|S|+1<\infty$. By 2.3, there exists a subfield $K_{1}$ of $K$ containing $K_{0}$ such that $vK_{1}/vK_{0}$ is torsion, $vK/vK_{1}$ is torsion-free, $Kv/K_{1}v$ is algebraic and purely inseparable and $v|_{K_{1}}$ is henselian. As the extension $Kv/K_{1}v$ is purely inseparable and ${\mathsf{char}(Kv)\neq 2}$, it is in particular a direct limit of normal extensions of odd degree. We conclude by 5.5 that $\widehat{u}(Kv)\leqslant\widehat{u}(K_{1}v)$. As the quotient group $vK/vK_{1}$ is torsion-free, we have $[vK_{1}:2vK_{1}]\leqslant[vK:2vK]$. Since $vK_{1}/vK_{0}$ is torsion, we conclude that $\mathsf{rk}(v|_{K_{1}})=\mathsf{rk}(v|_{K_{0}})<\infty$. Therefore, in view of the previous part, we have $\widehat{u}(K_{1})=[vK_{1}:2vK_{1}]\cdot\widehat{u}(K_{1}v)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. Letting $F_{1}=K_{1}(S)$, we have that $F_{1}/K_{1}$ is a function field in one variable and obtain that $\mathsf{dim}(\varphi)\leqslant u(F_{1})\leqslant 2\widehat{u}(K_{1})\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$. Having this for any anisotropic torsion form $\varphi$ defined over any function field in one variable $F/K$, we conclude that $\widehat{u}(K)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. Since the other inequality also holds, we have established the equality in all generality. ∎

### v2

@@ 5. The generalised strong $u$-invariant
We are now prepared to study the $u$-invariant of function fields in one variable. Let $K$ be a field. Extending [4, Sect. 5] (by assuming nothing further on $K$), we define
@@
We call $\widehat{u}(K)$ the generalised strong $u$-invariant of $K$. For nonreal fields of characteristic different from $2$, $\widehat{u}$ coincides with the ‘strong $u$-invariant’ defined in  [17].
@@
Let $L/K$ be a finite field extension. Then
$u(L)\leqslant\hbox{$\frac{1}{2}$}u(L(X))\leqslant\widehat{u}(K)\,.$
If $L/K$ is a simple extension or $\mathsf{char}(K)\neq 2$, then $u(L)\leqslant\hbox{$\frac{1}{2}$}u(K(X))\leqslant\widehat{u}(K)$.
@@
Since $L$ is the residue field of a discrete valuation on $L(X)$ and since $L(X)$ is real if and only if $L$ is real, the first inequality follows from $4.$ The second inequality is clear from the definition of $\widehat{u}(K)$. The refinement under the assumption that $L/K$ is simple or $\mathsf{char}(K)\neq 2$ now follows by [5, Proposition 6.1] and its proof. ∎
@@
Assume that $u(L)=\widehat{u}(K)$ for every finite field extension $L/K(\sqrt{-1})$. Then $u(F)=2\,\widehat{u}(K)$ for every function field in one variable $F/K$.
@@
Consider a function field in one variable $F/K$. By the definition of $\widehat{u}(K)$, we have $u(F)\leqslant 2\widehat{u}(K)$. We fix $x\in F$ transcendental over $K$. There exists a discrete valuation $w$ on $F$ which is trivial on $K$ and with $w(x^{2}+1)>0$. Then $Fw$ is a finite field extension of $K(\sqrt{-1})$, whence $u(Fw)=\widehat{u}(K)$ by the hypothesis. Since $[wF:2wF]=2$, we conclude by 4.4 that $u(F)\geqslant 2\widehat{u}(K)$. Therefore $u(F)=2\widehat{u}(K)$. ∎
The following is a partial extension of [13, Theorem] with no assumption on $\mathsf{char}(K)$. It is also a variation of [2, Theorem 7.3].
@@
The following are equivalent:
  1. $(i)$
    $K(\sqrt{-1})$ has no finite field extension of even degree,
  2. $(ii)$
    $u(F)=2$ for every function field in one variable $F/K$,
  3. $(iii)$
    $u(F)<4$ for some function field in one variable $F/K$.
If these equivalent conditions hold, then $\widehat{u}(K)=1$, otherwise $\widehat{u}(K)\geqslant 2$.
Proof.
[…]
Let $F/K$ be an arbitrary function field in one variable. Assume that $K(\sqrt{-1})$ has a finite extension of even degree. Then the proof of [2, Theorem 7.3] shows that there exists an anisotropic totally indefinite $2$-fold Pfister form $\pi$ over $F$. Then $\pi$ has signature zero at every ordering of $F$. Using Pfister’s Local-Global Principle [21, Chap. VIII, Theorem 6.14] if $K$ is real, it follows that $\pi$ is a torsion form. We conclude that $u(F)\geqslant\mathsf{dim}(\pi)\geqslant 4$. ∎
Note that by its definition, $\widehat{u}(K)$ is always in $\frac{1}{2}\mathbb{N}\cup\{\infty\}$, but we do not know whether it is always in $\mathbb{N}$. On the other hand, we do not know of any example where $\widehat{u}(K)$ is not a power of $2$.
In the following example and the subsequent statement, we use the concept of $\mathcal{C}_{i}$-fields (where $i\in\mathbb{N}$) introduced by S. Lang in [22].
Example 5.5.
Let $i\in\mathbb{N}$ be such that $K$ is a $\mathcal{C}_{i}$-field. Then $u(K)\leqslant 2^{i}$. Moreover, for any $j\in\mathbb{N}$ and any field extension $L/K$ of transcendence degree $j$, [31, Theorem 2a] (which is [22, Theorem 6] without the hypothesis on ‘normic forms’) yields that $L$ is a $\mathcal{C}_{i+j}$-field, whereby $u(L)\leqslant 2^{i+j}$. Applying this with $j=1$, we obtain that $\widehat{u}(K)\leqslant 2^{i}$.
Theorem 5.6 (Lang).
Let $(K,v)$ be a henselian discretely valued field with $\mathsf{char}(K)=0$ and algebraically closed residue field $Kv$. Then $K$ is a $\mathcal{C}_{1}$-field. In particular, for any $i\in\mathbb{N}$, we have $u(F)\leqslant 2^{i+1}$ for any field extension $F/K$ of transcendence degree $i$. Furthermore $\widehat{u}(K)=2$.
Proof.
That $K$ is a $\mathcal{C}_{1}$-field is proven in [22, Theorem 10] in the case where $(K,v)$ is complete. Let $K^{v}$ denote the completion of $K$ with respect to $v$. By [20, Theorem 5.9], the hypotheses on $(K,v)$ imply that $K$ is existentially closed in $K^{v}$. In other terms, every multivariate polynomial over $K$ has a zero over $K$ provided that it has a zero over $K^{v}$. Since $K^{v}$ is a $\mathcal{C}_{1}$-field, we conclude that $K$ is a $\mathcal{C}_{1}$-field.
As $K$ carries a discrete valuation, $K(\sqrt{-1})$ has a quadratic field extension, so 5.4 yields that $\widehat{u}(K)\geqslant 2$. The rest follows by $5.$ ∎
Proposition 5.7.
Let $v$ be a valuation on $K$. Then
@@ 3. $(iii)$
Consider an arbitrary function field in one variable $E/Kv$. By [3, Lemma 5.6], there exists a function field in one variable $F/K$ and an extension $w$ of $v$ to $F$ such that $Fw\simeq E$, and if $E$ is nonreal, then so is $F$. By 4.4, we obtain that $u(F)\geqslant[wF:2wF]\cdot u(Fw)$. By 2.2, since Fw=E{Fw=E} and $E/Kv$ is transcendental, the quotient group $wF/vK$ is finite. Hence 2.1 yields that $[wF:2wF]=[vK:2vK]$. Therefore
$vK:2vK]\cdot u(E)=[wF:2wF]\cdot u(Fw)\leqslant u(F)\leqslant 2\widehat{u}(K)\,.$
@@
In 6.2 we will see that the inequality in 5.7 becomes an equality when $v$ is henselian.
Proposition 5.8.
Let $v$ be a valuation on $K$ and let $F/K$ be a function field in one variable. Then there exists a function field in one variable $E/Kv$ such that $u(F)\geqslant 2\cdot[vK:2vK]\cdot\lfloor\frac{u(E)}{2}\rfloor$, and if $F$ is nonreal, then $u(F)\geqslant[vK:2vK]\cdot u(E)$.
Proof.
We choose a valuation $w$ on $F$ extending $v$ such that $Fw/Kv$ is a function field in one variable. (This can be done by fixing $x\in F$ transcendental over $K$, taking the Gauss extension of $v$ to $K(x)$ with respect to $x$ and then extending this valuation further to $F$.) By 2.2, the quotient group $wF/vK$ is finite. Therefore we have $[wF:2wF]=[vK:2vK]$, by 2.1. We now conclude by 4.4. ∎
@@
By 5.1, we have $u(K^{\prime})\leqslant\widehat{u}(K^{\prime})$. To show that $\widehat{u}(K^{\prime})\leqslant\widehat{u}(K)$, consider a function field in one variable $F^{\prime}/K^{\prime}$ and an anisotropic torsion form $\varphi$ over $F^{\prime}$. There exists a finitely generated subextension $F/K$ of $F^{\prime}/K$ such that $\varphi$ is defined over $F$ and is a torsion form over $F$. Moreover, we may choose $F$ to be transcendental over $K$. Then $F/K$ is a function field in one variable and $\varphi$ is an anisotropic torsion form over $F$. Hence $\mathsf{dim}(\varphi)\leqslant u(F)\leqslant 2\widehat{u}(K)$. Since this holds for any anisotropic torsion form $\varphi$ over $F^{\prime}$, it follows that $u(F^{\prime})\leqslant 2\widehat{u}(K)$. Having this for any function field in one variable $F^{\prime}/K^{\prime}$, we conclude that $\widehat{u}(K^{\prime})\leqslant\widehat{u}(K)$.
Assume now that $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree. It remains to show that in this case $\widehat{u}(K)\leqslant\widehat{u}(K^{\prime})$. Consider a function field in one variable $F/K$ and an anisotropic torsion form $\varphi$ over $F$. There exists a function field in one variable $F^{\prime}/K^{\prime}$ such that $F\subseteq F^{\prime}$ and $F^{\prime}=K^{\prime}F$. Since $K^{\prime}/K$ is a direct limit of finite normal extensions of odd degree, so is $F^{\prime}/F$. Therefore, Springer’s Theorem for odd degree extensions [21, Chap. VII, Theorem 2.7] yields that $\varphi$ remains anisotropic over $F^{\prime}$. Since it also remains a torsion form over $F^{\prime}$, we obtain that $\mathsf{dim}(\varphi)\leqslant u(F^{\prime})\leqslant 2\widehat{u}(K^{\prime})$. Since this holds for all anisotropic torsion forms $\varphi$ over $F$, we see that $u(F)\leqslant 2\widehat{u}(K^{\prime})$. Having this for all function fields in one variable $F/K$, we conclude that $\widehat{u}(K)\leqslant\widehat{u}(K^{\prime})$. ∎
For a general algebraic extension $K^{\prime}/K$, a strict inequality $\widehat{u}(K^{\prime})<\widehat{u}(K)$ may occur. For example, if $K^{\prime}$ is an algebraic closure of $K$ while $K(\sqrt{-1})$ has a finite field extension of even degree, then 5.4 yields that $\widehat{u}(K^{\prime})=1<\widehat{u}(K)$.
Question 5.10.
@@
Assume that $\mathsf{char}(K)\neq 2$ and $u(K(X,Y))=4u(K)$. Then $\widehat{u}(K)=u(K)$.
@@
By 5.9, we have $u(K)\leqslant\widehat{u}(K)$. To show the opposite inequality, consider an arbitrary function field in one variable $F/K$. Then $F$ is isomorphic to a finite field extension of $K(X)$, so that 5.1 and the hypothesis yield that $u(F)\leqslant\frac{1}{2}u(K(X,Y))=2u(K)$. This shows that $\widehat{u}(K)\leqslant u(K)$. ∎
The inequality $u(K(X))\geqslant 2u(K)$ has an analogue for the strong $u$-invariant.
@@
Consider a function field in one variable $F/K$. Then $F(X)/K(X)$ is a function field in one variable and $u(F(X))\geqslant 2u(F)$, by 5.1. Therefore we have $2\widehat{u}(K(X))\geqslant u(F(X))\geqslant 2u(F)$. Having this for every function field in one variable $F/K$ yields the claimed inequality. ∎
Question 5.13.
@@
Let $K$ denote the quadratic closure of $\mathbb{Q}$. Choose an arbitrary proper finite field extension $L/K$ (e.g. $L=K(\sqrt[5]{6})$). Then $L$ is nonreal, and by [21, Chap. VII, Cor. 7.11] we have $|{L}^{\times}/{L}^{\times 2}|=\infty$. In particular $u(L)\geqslant 2$. It follows by 5.1 that $u(K(X))\geqslant 2u(L)\geqslant 4>2=2u(K)$.
@@ 6. Function fields in one variable over a henselian valued field
We will now prove our main result and subsequently present various situations where it can be applied. We need a last step of preparation. For a valuation $w$ on a field $F$, we denote by $F^{w}$ the corresponding completion. Let $K$ be a field with $\mathsf{char}(K)\neq 2$.
@@
Let $v$ be a henselian rank-$1$ valuation on $K$ and $\mathsf{char}(Kv)\neq 2$. Let F/K{F/K} be a function field in one variable and $w$ a rank-$1$ valuation on $F$ with $\mathcal{O}_{v}\subseteq\mathcal{O}_{w}$. Then
@@
Since $2\in{\mathcal{O}}^{\times}_{v}\subseteq{\mathcal{O}}^{\times}_{w}$, we have that $\mathsf{char}(Fw)\neq 2$. Note that $w$ extends naturally to a complete valuation $w^{\prime}$ on $F^{w}$ with $w^{\prime}F^{w}=wF$ and $F^{w}w^{\prime}=Fw$. Hence $w^{\prime}$ is a complete rank-$1$ valuation on $F$ and $\mathsf{char}(F^{w}w^{\prime})\neq 2$. It follows by [15, Theorem 1.3.1 (Hensel’s Lemma)] that $1+\mathfrak{m}_{w^{\prime}}\subseteq{(F^{w})}^{\times 2}$. The equality on the left follows now from 4.6 applied to $(F^{w},w^{\prime})$. It remains to show that $wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$.
@@
$(1)$ Suppose that $Fw/Kv$ is a finite field extension and $wF/vK$ is a finitely generated $\mathbb{Z}$-module with $\mathsf{rrk}(Fw/Kv)=1$. It follows by 2.1 that $wF:2wF]=2\cdot[vK:2vK]$. By 5.9 we have that $u(Fw)\leqslant\widehat{u}(Kv)$. From this the claimed inequality follows.
$(2)$ Suppose that $Fw/Kv$ is a function field in one variable and $wF/wK$ is finite. Then by 2.1, $[wF:2wF]=[vK:2vK]$, and $u(Fw)\leqslant 2\widehat{u}(Kv)$, which implies the claimed inequality.
$(3)$ Suppose that $Fw/Kv$ is an algebraic field extension and $wF/vK$ is a torsion group. Then $u(Fw)\leqslant\widehat{u}(Kv)$ and $wF:2wF]\leqslant[vK:2vK]$ by 2.1. In this case we obtain even that $wF:2wF]\cdot u(Fw)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$.
$(4)$ Assume finally that $w|_{K}$ is trivial. In this case $w$ is a discrete valuation, so we may assume that $wF=\mathbb{Z}$, and the residue field $Fw$ is a finite extension of $K$. Hence $v$ extends to a valuation $v^{\prime}$ on $Fw$, and as $v$ is henselian, so is $v^{\prime}$. It follows by 4.6 that $u(F^{w})=[wF:2wF]\cdot u(Fw)=2u(Fw)$ and $u(Fw)=[v^{\prime}(Fw):2v^{\prime}(Fw)]\cdot u((Fw)v^{\prime})$. Since the extension $(Fw)v^{\prime}/Kv$ is finite, we have that $u((Fw)v^{\prime})\leqslant\widehat{u}(Kv)$, by 5.9. Furthermore, $v^{\prime}(Fw):vK]\leqslant[Fw:K]<\infty$, whence $v^{\prime}(Fw):2v^{\prime}(Fw)]=[vK:2vK]$ by 2.1. Thus $u(Fw)=[vK:2vK]\cdot u((Fw)v^{\prime})\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. ∎
We come to our generalisation of [36, Theorem 3] and [4, Cor. 6.4]. The proof crucially depends on the local-global principle for isotropy from [28, Cor. 3.9], or more precisely, on the slightly extended version provided in [3, Theorem 4.4], where the valuation on the base field is assumed of rank $1$ and henselian, but not necessarily complete.
@@
Furthermore, if $u(E)=2\widehat{u}(Kv)$ for every function field in one variable $E/Kv$, then $u(F)=2[vK:2vK]\cdot\widehat{u}(Kv)$ for every function field in one variable $F/K$.
@@
By 5.7 we have that $\widehat{u}(K)\geqslant[vK:2vK]\cdot\widehat{u}(Kv)$. We now show the converse inequality.
@@
In the next step, we assume that $v$ has arbitrary finite rank $n$ and show the statement in this situation. If $n=0$, then $v$ is trivial, whereby $Kv=K$, $[vK:2vK]=1$, and hence trivially $\widehat{u}(K)=\widehat{u}(Kv)=[vK:2vK]\cdot\widehat{u}(Kv)$. The case where $n=1$ is already established, and we will use it in the induction step. Assume that $n>0$. Then the value group $vK$ contains a unique maximal proper convex subgroup $\Delta$. It follows that the quotient group $vK/\Delta$ is a torsion-free group of rank $1$, $\Delta$ has rank n1n-1, and the ordering on $vK$ induces an ordering on $\Gamma=vK/\Delta$. Mapping $x\in K$ to $v(x)+\Delta$ defines a valuation $v^{\prime}$ on $K$ of rank $1$ with $v^{\prime}K=\Gamma$. The residue field $Kv^{\prime}$ of $v^{\prime}$ carries a valuation $\overline{v}$ with value group $\Delta$ and hence of rank n1n-1. Furthermore, $\mathcal{O}_{v}\subseteq\mathcal{O}_{v^{\prime}}$, and $\mathfrak{m}_{v^{\prime}}$ is an ideal of $\mathcal{O}_{v}$ such that $\mathcal{O}_{v}/\mathfrak{m}_{v^{\prime}}\simeq\mathcal{O}_{\overline{v}}$. Since $v$ is henselian, the valuations $v^{\prime}$ on $K$ and $\overline{v}$ and $Kv^{\prime}$ are both henselian. By the induction hypothesis, we obtain that $\widehat{u}(Kv^{\prime})=[\overline{v}(Kv^{\prime}):2\overline{v}K^{\prime}(Kv^{\prime})]\cdot\widehat{u}((Kv^{\prime})\overline{v})$, and by the case where $n=1$, we have that $\widehat{u}(K)=[v^{\prime}K:2v^{\prime}K]\cdot\widehat{u}({Kv^{\prime}})$. Furthermore, $(Kv^{\prime})\overline{v}=Kv$, $v^{\prime}K=\Gamma$ and $\overline{v}(Kv^{\prime})=\Delta$. Therefore $\widehat{u}(K)=[\Gamma:2\Gamma]\cdot\widehat{u}(Kv^{\prime})=[\Delta:2\Delta]\cdot[\Gamma:2\Gamma]\cdot\widehat{u}(Kv)$. Finally, since $vK$ and $vK/\Delta$ are torsion-free, it readily follows that $vK:2vK]=[\Gamma:2\Gamma]\cdot[\Delta:2\Delta]$. We conclude that $\widehat{u}(K)=[vK:2vK]\cdot\widehat{u}(Kv)$, as desired.
@@
We set $K_{0}=\mathbb{F}(S\cap K)$. Since the absolute transcendence degree of $K_{0}$ is bounded by $|S|$, it follows by [15, Theorem 3.4.3] that $\mathsf{rk}(v|_{K_{0}})\leqslant|S|+1<\infty$. By 2.3, there exists a subfield $K_{1}$ of $K$ containing $K_{0}$ such that $vK_{1}/vK_{0}$ is torsion, $vK/vK_{1}$ has no elements of order $2$, $Kv/K_{1}v$ is algebraic and purely inseparable and $v|_{K_{1}}$ is henselian. As the extension $Kv/K_{1}v$ is purely inseparable and $\mathsf{char}(Kv)\neq 2$}, it is in particular a direct limit of normal extensions of odd degree. We conclude by 5.9 that $\widehat{u}(Kv)\leqslant\widehat{u}(K_{1}v)$. As $vK/vK_{1}$ has no elements of order $2$, we have $vK_{1}:2vK_{1}]\leqslant[vK:2vK]$. Since $vK_{1}/vK_{0}$ is torsion, we conclude that $\mathsf{rk}(v|_{K_{1}})=\mathsf{rk}(v|_{K_{0}})<\infty$. Therefore, in view of the previous part, we have $\widehat{u}(K_{1})=[vK_{1}:2vK_{1}]\cdot\widehat{u}(K_{1}v)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. Letting $F_{1}=K_{1}(S)$, we have that $F_{1}/K_{1}$ is a function field in one variable and obtain that $\mathsf{dim}(\varphi)\leqslant u(F_{1})\leqslant 2\widehat{u}(K_{1})\leqslant[vK:2vK]\cdot 2\widehat{u}(Kv)$. Having this for any anisotropic torsion form $\varphi$ defined over any function field in one variable $F/K$, we conclude that $\widehat{u}(K)\leqslant[vK:2vK]\cdot\widehat{u}(Kv)$. Hence, we have established the equality $\widehat{u}(K)=[vK:2vK]\cdot\widehat{u}(Kv)$ in all generality.
For the last part of the statement, assume now that, for every function field in one variable $E/Kv$ we have $u(E)=2\widehat{u}(Kv)$, and hence also $u(E)=u(Kv(T))$. Consider an arbitrary function field in one variable $F/K$. If $K$ is real, then it follows by 3.10 $(a)$ that $Kv$ is real, whereby $Kv(T)$ is real and $u(Kv(T))$ is even. Therefore, 5.8 yields that there exists a function field in one variable $E/Kv$ such that $u(F)\geqslant[vK:2vK]\cdot u(E)$. Since $u(E)=2\widehat{u}(Kv)$ and we already established that $u(F)\leqslant 2\widehat{u}(K)=2[vK:2vK]\cdot\widehat{u}(Kv)$, we conclude that $u(F)=2[vK:2vK]\cdot\widehat{u}(Kv)$. ∎